OpenStudy (anonymous):

The theoretical probability of undesirable side effects resulting from taking Grebex is 3 in 11. If 110 people take Grebex to lower their blood pressure how many will encounter undesirable side effects?

OpenStudy (anonymous):

\[\frac{3}{11}=\frac{x}{110}\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

so would it 33/110

OpenStudy (anonymous):

no

OpenStudy (anonymous):

or 30/110

OpenStudy (anonymous):

\[\frac{3}{11}=\frac{x}{110}\iff x=\frac{110\times 3}{11}\]

OpenStudy (anonymous):

It's 24 if rounded.. Right?

OpenStudy (anonymous):

you would expect the number of people with side effects to be larger than one

OpenStudy (anonymous):

x is 110/14

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Wrong onee, I'm sorry! I'm in the wrong open study. lol

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

330/11 is that right

OpenStudy (anonymous):

3 in 11 is the same as 3x in 11x where 3x is # with unde. side effects. Equation: 3x+11x=110 14x=110 x=110/14 # of 110=3x=3(110/14)=23.57,. round up to 24 (: I got it on this website. http://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.408582.html

OpenStudy (anonymous):

Does that help?